OpenStudy (anonymous):

find the GCF of 16a4b2 + 20a2b2 - 4ab2. please help? :C

OpenStudy (anonymous):

\[16a^4b^2=2 \times 2 \times 2 \times 2 \times a \times a \times a \times a \times b \times b\]\[20a^2b^2 = 2 \times 2 \times 5 \times a \times a \times b \times b\]\[4ab^2 = 2 \times 2 \times a \times b \times b\]how many a and b and 2 do they all have in common?